# A Theory of Recommendations

### v2

@@ 5.3.1 Preference Heterogeneity
Suppose Assumption 1 holds.
  1. 1.
    Suppose $\sigma<1$. Then, a mean-preserving spread of the type distribution increases the value of the recommendation system for $R>1/2$ and decreases it for $R<1/2$.
  2. 2.
    Suppose $\sigma>1$. Then, a mean-preserving spread of the type distribution decreases the value of the recommendation system for $R>1/2$ and increases it for $R<1/2$.
To gain some intuition, suppose $\sigma<1$ and $R>1/2$. Figure 1 shows that with $R>1/2$, only extreme types give buy recommendations for controversial products. [^30] Conversely, for $R<1/2$ only relatively extreme types give don’t-buy recommendations for controversial products. This can be seen by flipping the colors in Figure 1 so that it corresponds to $R=1/4$. Increasing $\beta$ corresponds to a mean-preserving spread of $F$ given $R$. Consequently, there are more such extreme senders and, hence, buy recommendations are increasingly likely to result from controversial products. This increases the probability of buy recommendations while lowering their objective value. The resolution of this trade-off, and thus whether the value of the recommendation system goes up or down, is determined by the odds ratio $\sigma$ (see Proposition 5). For instance, if $\sigma<1$, good products are rare compared to bad ones, so the value of the recommendation system is increasing in $\beta$. Therefore, the positive effect of revealing bad products by means of don’t-buy recommendations outweighs the negative effect of a reduced value conditional on receiving a buy recommendation. Thus, given a high recommendation threshold ($R>1/2$), a more polarized population benefits more from the recommendation system, as this means only bad products generate don’t-buy recommendations. Conversely, when the recommendation threshold is relatively low ($R<1/2$), a less polarized population centered around the middle gives don’t-buy recommendations only to bad products.
Figure 3: Population polarization
[…]
5.3.2 Prevalence of Controversial Products
Without controversial products, preference heterogeneity would not matter: All receivers would accept all recommendations, as they would reveal whether the product is good or bad. Hence, the value of the recommendation system would be independent of the preference distribution. More broadly, we now ask how an increasing prevalence of controversial products (an increase in $Q$) affects the value of a recommendation system. A natural conjecture is that recommendations become less valuable when controversial products are more likely. We find that this conjecture needs to be qualified.
Corollary 3.
Suppose Assumption 1 holds.
  • (i)
  • (ii)
Figure 4: Illustration of Corollary 3
(a) Panel A
(b) Panel B
(c) Panel C
Figure 4 illustrates the result. In the white area in Panel A, $V(\beta)$ is decreasing in $\beta$; in the shaded area, an interior maximizer exists. When good and bad products are similarly likely ($\sigma$ is close to $1$), a higher prevalence of controversial products indeed reduces $V(\beta)$. For very high or low values of $\sigma$, however, this need not be true. This surprising result reflects the multiplicative structure of $V(\beta)$. Under Assumption 1, $V(\beta)$ is the product of the buying probability $\pi^{B}$ and the objective value of the recommendation per buyer, $\Delta_{O}^{B}$. The effect of $Q$ on $\pi^{B}$ is negative only if $\sigma$ is sufficiently large (see Panel B). Intuitively, as $Q$ increases, good products are less likely, reducing the probability of buy recommendations. However, there are now more buy recommendations stemming from controversial products, as these have become more likely. $\beta$ and $\sigma$ determine which of these effects dominates. When $\beta$ is high, the greater likelihood that controversial products receive buy recommendations increases the total probability of buy recommendations, strengthening the positive effect. When $\sigma$ is high, the lower share of good products dominates, strengthening the negative effect. Similarly, one can show that the effect of $Q$ on $\Delta_{O}^{B}$ becomes negative only for small $\sigma$ (see Panel C). Again, this reflects countervailing forces. The negative effect results from the increased probability that a product with a buy recommendation is controversial. The positive effect reflects a change in the value of the outside option. The values of $\sigma$ and $\beta$ determine which effect dominates. [^31] When $\beta$ is small, the negative effect is weak, as only few types give buy recommendations for controversial products. When $\sigma$ is small, the outside option becomes more valuable, as bad products become less likely, weakening the positive effect.
6 Extensions
@@
Our benchmark model relies on several assumptions. First, the preference distribution of senders and receivers is identical. Second, there are only two levels of recommendations (“buy” and “don’t-buy”). Third, each receiver only has access to a single recommendation. We now relax each of these assumptions in turn.
@@ 6.1 Distinct Sender and Receiver Distributions
We modify the benchmark model by assuming that the sender and the receiver are drawn from different distributions $F$ and $G$, respectively. For instance, this enables us to take into account that the “early adopters” of a new product may have different preferences than the general population (similar to Acemoglu et al. (2022)). We show that the results of the benchmark case easily generalize to this setting. More importantly, the general setting allows us to treat the choice of the sender distribution as a potentially interesting design variable.
The sender distribution $F$ fully determines the probabilities of buy and don’t-buy recommendations, the posterior probabilities, and thus the objective and subjective effects and the expected payoffs, $V_{A}$ and $V_{N}$. Hence, all results on receiver behavior in Section 4 apply verbatim if $F$ is taken to be the sender distribution, as the receiver’s behavior is independent of the receiver distribution. The value of the recommendation system, however, depends on the receiver distribution $G$ as well, as it involves integration over this distribution. The next result generalizes Proposition 3.323232The notation for the quantities in Proposition 9 reflects their dependence on $F$ and $G$, respectively.
@@
Consider the model with distinct preference distributions of senders and receivers. Suppose the sender distribution $F$ satisfies Assumption 1. Then, all types accept all recommendations. Further, $V(R)=\pi^{B}_{F}[\Delta_{O,F}^{B}-\Delta_{S,F}^{B}\mathbb{E}_{G}[i]]$ is
  1. (i)
  2. (ii)
  3. (iii)
Recall that in the benchmark model symmetry of $F$ implies that all receivers accept all recommendations and the subjective effect of the recommendation cancels out. This is still true if $F$ is distinct from $G$ as long as $G$ is also symmetric and hence $\mathbb{E}_{G}[i]=0$. For more general distributions $G$, all receivers still accept all recommendations for the same reason as above. However, as the receiver distribution is not symmetric, the subjective effect does not cancel out anymore. Even so, the result qualitatively mirrors Proposition 5, as changing the threshold still leads to the same trade-off: Reducing the threshold increases the probability of a buy recommendation but decreases its value conditional on that recommendation. However, in the latter effect, the value is no longer purely objective but consists of the objective part $\Delta_{O}^{B}$ and a subjective part given by $-\Delta_{S}^{B}\mathbb{E}_{G}[i]$, reflecting the receiver distribution $G$. Further, as the subjective effect now influences the value of the system, so does the prevalence of controversial products by affecting the tipping point at which the designer opts for a high or a low threshold. In any case, when $\mathbb{E}_{G}[i]=0$, we obtain the result from the baseline model as a special case. Similarly, the other results in the benchmark model remain (qualitatively) valid. With asymmetrically distributed populations, Propositions 6 and 7 hold for arbitrary receiver distributions as long as we impose Assumption 2 on the sender distribution.
Allowing for differences in sender and receiver distributions makes it possible to address another important design issue. In some cases, it is natural to assume that a designer, for instance a platform, can influence the sender distribution or at least the receiver’s perception of it. Assuming that the platform has better knowledge about sender types, it may choose to share some of this information with the receivers or even restrict the option of sending recommendations to certain types of senders (and inform the receivers accordingly). For instance, many platforms such as Amazon inform browsing customers that “customers who bought this item also bought” another item.
There are different ways to model this. First, suppose that $F$ is symmetric and that the designer can render the sender distribution more or less precise using mean-preserving contractions or mean-preserving spreads, respectively. Intuitively, one would expect more precise knowledge about the sender to increase the value of the recommendation system. However, the logic of Proposition 9 shows that this is not necessarily the case. As in Section 5.3.1, the effects of mean-preserving contractions and spreads can be captured as changes in $\beta=F(1/2-R)$, just like changes in $R$. With mild adaptions, the proof of Proposition 9 can thus be used to show that whether more or less precise information about the sender increases the value of recommendations depends on the parameters.
Second, suppose that the sender and receiver distributions are fixed but that the designer can provide the receivers with a signal about the sender that allows for inference about the sender’s type. To make the point as starkly as possible, suppose the signal fully reveals the sender’s type. If the sender is revealed to be type $i=0$, all types accept all recommendations, as both types of recommendations contain only objective information. In contrast to Proposition 2—where this behavior is shown to ensue with extreme thresholds—this would arise for all thresholds $R\in(0,1)$. Intuitively, as the type $i=0$ is indifferent between controversial products, her recommendation never reveals subjective information. In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too. [^33] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
Finally, suppose that the sender and receiver distributions are not independent but somehow correlated. Recommendations would then be “tailored” to the receiver $\bar{i}$, as they would be drawn from a sender distribution $F(i|\bar{i})$, resulting in type-specific posteriors $p^{r}_{j}(i)$ for $j\in\{H,1,2,L\}$. In such a setting, an analogous result to Proposition 1 still holds, establishing that either all all or no recommendations are accepted. Further analysis would give rise to similar complications as in Section 5.2.2 with asymmetric distributions, as the the correlation between sender and receiver distribution would affect the set of types who accept all recommendations.
6.2 Multiple Recommendation Thresholds
We now analyze the effects of reducing the coarseness of the recommendation system. We maintain all assumptions of the benchmark case, except that there are two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$ and three types of recommendations $r\in\{D,N,B\}$ which we interpret as “don’t-buy,” “neither buy nor don’t-buy,” and “buy.” [^34] Such a two-threshold structure is found to be optimal in models of strategic communication in the context of social learning by Smirnov and Starkov (2024) and Bénabou and Vellodi (2024). The sender gives a buy recommendation if the payoff is at least $R_{2}$, a don’t-buy recommendation if it is below $R_{1}$, and neither if it is in $[R_{1},R_{2})$. We maintain Assumption 1.
The proof of Proposition 10 in Appendix A.3 provides the receiver’s posteriors. For buy and don’t-buy recommendations, these posteriors are analogous to the benchmark, except that the different thresholds in each case have to be taken into account. As in the benchmark model, the highest and lowest recommendations rule out the bad and good product versions, respectively. The intermediate recommendation $r=N$, however, rules out both of these versions, so the receiver knows that the recommended product is controversial.
We say that a receiver accepts all recommendations if he buys if $r=B$ and does not buy if $r=D$, allowing for either behavior if $r=N$. Assumption 1 again implies that the receiver always buys a product with $r=B$ and never buys if $r=D$. As noted above, $r=N$ rules out both uncontroversial products. Given the symmetric population, $r=N$ leaves the relative likelihood of the controversial products unchanged. Hence, whether a type $i$ wants to buy a product with $r=N$ depends on (i) how likely a good product is and (ii) how much he (dis-)likes the more likely controversial product. As in the benchmark model, we can derive the indifferent type so that, depending on parameters, either all, only sufficiently aligned, or no types buy the product with an intermediate recommendation.
$V(R)$ now consists of two components. First, as in the benchmark, all types accept $r=B$, resulting in the component $\pi^{B}\Delta_{O}^{B}$. [^35] As before, accepting don’t-buy recommendations does not create value. Moreover, $r=N$ can create value for the types that buy following that neutral recommendation. As noted above, parameter values determine which types do this. Despite this ambiguity, we can formulate the following result.
Proposition 10.
@@ 6.2 Multiple Recommendation Thresholds
Consider the model with multiple recommendation thresholds. Suppose Assumption 1 holds. $V(R)$ is (i) decreasing in $R_{1}$ if $\sigma<1$ and (ii) increasing in $R_{2}$ if $\sigma>1$.
@@
The result provides an incomplete characterization of the optimal recommendation system: It establishes that one of the two thresholds should be extreme so that one recommendation essentially becomes objective. Thus, as in Proposition 5, the designer maximizes the probability of an objective recommendation, considering which uncontroversial product version is more likely a priori. The optimal value of the threshold that is not pinned down by Proposition 10 depends on the model parameters unless $q_{1}=q_{2}$. In this degenerate case, any value for the other threshold is optimal, but it is also possible that both the high and the low recommendations are essentially turned objective. [^36] This arises for instance for $F(i)=1/2+i$ with parameters $q_{1}=0.7,\ q_{2}=0.1,$ and $q_{H}=0.03.$
@@ 6.3 Multiple Recommendations
We assume that receivers have access to a number $b$ of buy recommendations and a number $d$ of don’t-buy recommendations. As in the benchmark model, recommendations come from randomly drawn senders. We add the assumption that they are drawn independently. [^37] In Appendix A.4, we provide the posteriors, depending on $b$ and $d$. Just as before, a single buy recommendation is enough to rule out a bad product, and a single don’t-buy recommendation is enough to rule out a good product. Consequently, when the receiver obtains mixed recommendations, that is, both buy and don’t-buy recommendations, he rules out both uncontroversial product versions. [^38] In Sun (2011), a higher variance in the recommendations corresponds to a niche product that some consumers like and some don’t. Similarly, in our paper, mixed recommendations (and thus a higher variance) result from controversial products that, indeed, some consumers like more than others.
To make our main point that multiple recommendations need not increase the value of the recommendation system as starkly as possible, we focus on the case of “infinite learning” where the receiver has access to an unlimited number of recommendations.
@@
Consider the model with multiple recommendations. Suppose Assumption 1 holds. Let $\lambda:=q_{1}/q_{2}$. Compared with the optimal recommendation system in the benchmark model, infinite learning does not increase the value of the recommendation system if and only if either (i), (ii), or (iii) holds:
@@
To grasp the result intuitively, recall the receiver’s behavior in the single-recommendation system. With a threshold of $R\rightarrow 0$, the receiver buys the recommended product unless it is bad. For $R\rightarrow 1$, he does so only if it is good. In the benchmark model, V(R) is decreasing in $R$ if $\sigma<1$ and increasing if $\sigma>1$. Lemma 4 in Appendix A.4 describes optimal receiver behavior with infinite learning. It shows that behavior coincides with the optimally designed single-recommendation system if the product is uncontroversial: Good products are bought but bad ones are not. Thus, the infinite-learning system potentially differs from the benchmark only for controversial products, which generate mixed recommendations. When $\sigma$ is very low, bad products are relatively likely, so the outside option looks relatively bleak. A consumer facing a product with mixed recommendations will thus buy this (controversial) product. In contrast, when $\sigma$ is very high, any such consumer will buy the outside option, as it is likely to be good. Thus, when the probabilities of the uncontroversial product versions are not too similar, the behavior with a single recommendation and with multiple recommendations coincides, and multiple recommendations do not increase the value of the recommendation system. However, multiple recommendations do affect behavior if the probabilities of the good and bad products are similar. For instance, when $\lambda>1$, low types who value the less likely controversial product version $(0,1)$ more strongly than product version $(1,0)$ will opt for the outside option rather than the controversial product. Then, the system’s value with multiple recommendations is higher than that of the single-recommendation system, as it provides more granular information. Inspecting the conditions on $\sigma$ in (ii) and (iii), we find that the more extreme (further away from 1) the odds ratio of the controversial products is, the more extreme the odds ratio of the uncontroversial product versions needs to be for there to be no value in additional recommendations.

### v3

@@ 5.2 Demand-Maximizing Recommendation Systems
Suppose $q_{H}\geq 1/2$. Then, demand is maximized for $R\rightarrow 0$.
While this result holds without Assumption 1 or Assumption 2, it restricts the analysis to the case that high quality goods are quite common. This complication arises because, contrary to the symmetric case, raising $R$ does not simply reduce the probability of buy recommendations—it also changes who follows them: As $R$ increases, controversial products are less likely to be recommended, and the composition of receivers who accept recommendations shifts toward those whose preferences are more aligned with the sender.
To understand the intuition for Proposition 9, recall that demand can come from two sources: types who accept buy recommendations and types who reject don’t-buy recommendations and choose to buy the product anyway. This suggests two potential strategies for maximizing demand. One option is to set $R$ low, ensuring that many buy recommendations are issued and that all types follow them (see Proposition 2). Another option would be to set an intermediate $R$ where not all types accept recommendations, thus relying on many people rejecting don’t-buy recommendations so that demand remains high even when buy recommendations are rarer. However, when $q_{H}\geq 1/2$, buy recommendations are always more frequent than don’t-buy recommendations. Consequently, the second approach is always dominated by the first approach. Thus, the optimal strategy is to maximize the frequency of buy recommendations by setting $R\to 0$.
This result stands in stark contrast with value maximization under asymmetry. According to Proposition 8, choosing as low a threshold as possible is optimal if and only if $\sigma\leq\min\left\{\frac{1-Q(a+1)}{a-Q(a+1)},\frac{a-Q(a+1)}{1-Q(a+1)}\right\}$ holds. However, given $q_{H}\geq 1/2$, it can be shown that this condition is never satisfied. This means that the objectives of demand and value maximization are never aligned in this setting: while demand is always maximized at $R\to 0$, value maximization typically favors a higher threshold.
Intuitively, note that good products are relatively frequent when $q_{H}\geq 1/2$, meaning that buy recommendations are already quite likely. In this case, the emphasis in value maximization shifts from frequency to informativeness: Since buy recommendations are common, their informativeness matters more than their frequency. Increasing $R$ filters out more controversial products, making buy recommendations more selective and thus more valuable. In contrast, demand maximization focuses exclusively on making buy recommendations as frequent as possible.
As the requirement that $q_{H}\geq 1/2$ is quite strong, we carry out a complementary numerical analysis (see Appendix B) for the case where it does not hold. For this analysis, we need to invoke Assumption 2. The results suggest that demand is monotonically decreasing in $R$, just as in the symmetric case. Therefore, as in the case of a symmetric population, demand maximization will still be aligned with value maximization for some parameters but not for others.
6 Robustness and Further Design Considerations
@@
We now relax the assumptions that the preference distribution of senders and receivers is identical, that there are only two levels of recommendations (“buy” and “don’t-buy”) and that each receiver only has access to a single recommendation.
@@ 6.1 Distinct Sender and Receiver Distributions
Allowing for distinct sender and receiver distributions is interesting for two reasons. First, it captures situations such as when early adopters differ systematically from the broader population (similar to Acemoglu et al. (2022)). Second, and more importantly, it may be possible to shape the sender population or influence how receivers perceive it.
We first show that our earlier results extend naturally to the case of distinct sender and receiver distributions. Let the sender be drawn from $F$ and the receiver from $G$. If $F$ satisfies Assumption 1 or 2, the results on value- and demand-maximizing thresholds from Sections 4 and 5 remain valid. For example, Proposition A.2 in the appendix generalizes Proposition 5, showing that all receiver types accept recommendations under symmetry of $F$, and that $V(R)$ remains monotonic in $R$ with the slope depending on $\sigma$. Similarly, the potential conflict between value and demand objectives persists.
Beyond such robustness results, differing sender and receiver distributions give rise to new design considerations. In many cases, platforms may have partial control over the sender population or the information receivers have about it. For instance, Amazon’s phrasing “customers who bought this item also bought…” implicitly informs the receiver about who the senders are. We outline different ways to model this.
First, suppose the designer can manipulate the informativeness of a symmetric sender distribution $F$ through mean-preserving spreads or contractions. Though it may seem that making the sender distribution more precise would raise the value of the system, this is not always true. As in Section 4.3.1, changes in precision translate into changes in $\beta=F(1/2-R)$. With mild adaptations, the proof of Proposition A.2 can be used to show how the effect on value depends on the overall environment.
Second, suppose that the sender and receiver distributions are fixed, but that the designer can provide the receivers with a signal about the sender’s type. To make the point as starkly as possible, suppose that the signal fully reveals the type. If the sender is revealed to be $i=0$, all types accept recommendations, as the latter contain only objective information. In contrast to Proposition 2, where this behavior requires extreme thresholds, this would arise for all thresholds $R\in(0,1)$. Intuitively, as $i=0$ is indifferent between controversial products, her recommendation does not reveal subjective information. In contrast, if the signal reveals a sender of type $i\neq 0$, the receiver will be able to learn about controversial products, too. [^20] For instance, if $i=1$ and $R\in(0,1)$, a buy recommendation reveals that the product is either of version $(1,1)$ or $(1,0)$ Thus, having a sender who is partial to some controversial product is not necessarily a bad thing.
Third, suppose sender and receiver types are correlated so that recommendations are “tailored” to the receiver. In this case, the sender distribution becomes type-dependent, $F(i|\bar{i})$, and induces posterior beliefs that differ across receivers. While a suitably modified version of Proposition 1 still ensures that either all or no types follow recommendations, the set of types who do so now depends on the joint distribution. This leads to complications in further analyses similar to those discussed in Section 5 for the case of asymmetric populations.
6.2 Multiple Recommendation Thresholds
We now consider a recommendation system with finer granularity. Specifically, we relax the benchmark assumption of binary recommendations by introducing a third, intermediate option. Such richer recommendation structures appear in real-world platforms and are also shown to be optimal in models of strategic communication under social learning (e.g., Smirnov and Starkov, 2024; Bénabou and Vellodi, 2024).
Formally, we introduce two threshold levels $R_{1},R_{2}\in(0,1)$ with $R_{2}\geq R_{1}$, and allow for three recommendation types $r\in\{D,N,B\}$: “don’t-buy,” “neutral,” and “buy.” The sender issues a buy recommendation if her payoff is at least $R_{2}$, a don’t-buy if it is below $R_{1}$, and a neutral recommendation otherwise. We maintain Assumption 1 throughout.
As in the benchmark model, the highest and lowest recommendations convey objective information: $r=B$ rules out bad products, and $r=D$ rules out good ones. The neutral recommendation, by contrast, carries no objective content; it implies that the product is controversial, but does not distinguish between versions $(1,0)$ and $(0,1)$. This leads to the following result:
@@ 6.2 Multiple Recommendation Thresholds
Suppose there are recommendation thresholds $R_{2}\geq R_{1}$ and Assumption 1 holds. Then $V(\mathcal{R})$ is (i) decreasing in $R_{1}$ if $\sigma<1$, and (ii) increasing in $R_{2}$ if $\sigma>1$.
@@
This result mirrors the logic of Proposition 5: the designer maximizes the value of the recommendation system by ensuring that one of the thresholds produces recommendations that are essentially objective. The optimal threshold depends on which uncontroversial product version is more likely. If, for instance, bad products are more prevalent ($\sigma<1$), the low threshold $R_{1}$ should be set close to zero to ensure that don’t-buy recommendations are informative. The optimal value of the remaining threshold depends on the model parameters, unless $q_{1}=q_{2}$, in which case any choice yields the same outcome. In some cases, both thresholds may be set at extreme values, rendering both $r=B$ and $r=D$ essentially objective. [^21] This arises, for instance, under $F(i)=1/2+i$ with $q_{1}=0.7$, $q_{2}=0.1$, and $q_{H}=0.03$.
For a demand-maximizing platform, however, richer recommendation systems offer no benefit under Assumption 1. Since receivers of the neutral recommendation are only partially informed, they may or may not buy the product, depending on the environment. As shown in the proof of Proposition 10, this ambiguity leads to weakly lower demand than under a binary system with a very low threshold $R\to 0$, which maximizes the number of buy recommendations. Hence, the intermediate recommendation dampens demand without offsetting gains in platform revenue.
@@ 6.3 Multiple Recommendations
Formally, suppose the receiver observes $b$ buy and $d$ don’t-buy recommendations, each independently drawn from a population of senders as in the benchmark model. [^22] We derive the posteriors for general $(b,d)$ in Appendix A.5. A single buy recommendation is sufficient to rule out a bad product, and a single don’t-buy recommendation rules out a good product. Hence, when the receiver obtains a mix of both, he infers that the product is not uncontroversial—it must be one of the two controversial versions. [^23] See also Sun (2011), who relate recommendation variance to product “nicheness” in a similar spirit. To highlight the limits of informational gains from additional recommendations, we focus on the case of infinite learning, where the receiver observes an unbounded number of independent recommendations.
@@
Consider the model with multiple recommendations. Suppose Assumption 1 holds and let $\lambda:=q_{1}/q_{2}$. Then, infinite learning does not increase the value of the recommendation system relative to the optimally designed single-recommendation benchmark if and only if one of the following holds:
@@
The intuition is as follows. With infinitely many recommendations, the receiver can perfectly infer whether the product is good, bad, or controversial. As in the benchmark case, receivers buy good products and reject bad ones. Thus, any potential gain must come from improved decisions about controversial products. However, controversial products always generate mixed recommendations. Whether the receiver ultimately chooses to buy depends on his posterior belief and thus on the relative likelihood of the outside option being good or bad. If good products are relatively rare (that is, $\sigma$ is low), the outside option is unattractive, and the receiver may prefer a controversial product even if it is only an imperfect match. If good products are common (that is, $\sigma$ is high), the receiver is more willing to choose the outside option. In either extreme, the decision under infinite learning coincides with what could already be achieved by a single recommendation from an optimally calibrated recommendation system. Only when the uncontroversial versions are similarly likely and controversial products are sufficiently ambiguous does additional information add value.
Finally, note that providing more recommendations also does not help a demand-maximizing platform under Assumption 1. If the platform’s goal is to induce as many purchases as possible, it suffices to choose a single threshold $R\rightarrow 0$, which leads all senders to issue buy recommendations. Since mixed recommendations can only reduce the likelihood of purchasing controversial products, adding them weakly reduces demand.